# Near-Optimal Channel Estimation for Dense Array Systems

### v1

@@ IV-C Ice-Filling Versus Water-Filling
In this subsection, we put forward a more insightful interpretation to the ice-filling algorithm that establishes its connection with the water-filling principle. Generally speaking, the proposed ice-filling algorithm can be viewed as a quantization of the water-filling principle. Specifically, it is clear that the observation matrices generated by both the water-filling and the proposed ice-filling algorithms fall within the eigenspace of ${\bm{\Sigma}}_{\bf h}$. The major distinction is attributed to their “power allocation strategies”. The water-filling principle allocates $Q$ units of power to the $K$ eigenvectors using (8). On the other hand, the ice-filling algorithm transforms this power-allocation process into an eigenvector-assignment process, where the $K$ eigenvectors are repeatably assigned to the observation vectors in $Q$ timeslots. To be more precise, recall that $p_{k}$ in (8) denotes the optimal power allocated to the $k$-th eigenvector using the water-filling principle, satisfying $\sum_{k=1}^{K}p_{k}=Q$. For the sake of discussion, we introduce the definition of “pilot reuse frequency” as follows.
@@
Comparing the definitions of $p_{k}$ and $n_{k}$, we can intuitively interpret the $n_{k}$-timeslot reuse of the observation vector ${\mathbf{u}}_{k}$ for pilot transmission as allocating $n_{k}$ units of power to ${\mathbf{u}}_{k}$. More importantly, Definition 1 enables us to derive the analytical expressions for the eigenvalues $\{\lambda_{k}^{t}\}$ in Algorithm 2. Note that the eigenvalue update rule of the ice-filling algorithm can be rewritten as a recursive formula:
@@
where $\lambda_{k_{t}}^{t}$ is the largest eigenvalue from $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$. This recursive formula implies that whenever the $k$-th eigenvector is selected as an observation vector, the value of $\frac{\sigma^{2}}{\lambda_{k}^{t}}$ increases by 1, which naturally gives rise to the close-form expression of $\lambda_{k}^{t}$ in Lemma 2.
Lemma 2
The eigenvalue $\lambda_{k}^{t}$ and the pilot reuse frequency $n_{k}^{t}$ satisfy the relationship:
@@
Proof:
Equation (17) holds because the definition of $n_{k}^{t}$ suggests that the $k$-th eigenvector is selected by $n_{k}^{t}$ times during timeslots $1\sim t$, so that we have $\frac{\sigma^{2}}{\lambda_{k}^{t}}=n_{k}^{t}+\frac{\sigma^{2}}{\lambda_{k}}$. ∎
By comparing Lemma 2 with the optimal power allocation in (8), we can interpret our proposed ice-filling algorithm as a quantization/solidification version of the water-filling algorithm, which is elaborated below.
Interpretation of ice-filling: As depicted in Fig. 4, the water-filling principle allocates $Q$ units of water (power) to a vessel with $K$ channels, each having a unique base level $\frac{\sigma^{2}}{\lambda_{k}}$, $k\in\{1,2,\cdots,K\}$. By controlling the uniform water level $\beta$, the optimal power $p_{k}$ is determined by the gap between the base level and the water level. In contrast, the ice-filling algorithm transforms the total $Q$ units of water into $Q$ ice blocks, each containing one unit of power and being used for one pilot transmission. Our algorithm starts from an empty vessel and fills one ice block onto the channel having the deepest base surface $\frac{\sigma^{2}}{\lambda_{k_{0}}}$. This operation is equivalent to finding the largest eigenvalue $\lambda_{k_{0}}$ from $\{\lambda_{1},\cdots,\lambda_{K}\}$. Then, the eigenvector ${\mathbf{u}}_{k_{0}}$ is assigned to the first observation vector ${\mathbf{w}}_{1}$, and ice level of this channel increases from $\frac{\sigma^{2}}{\lambda_{k_{0}}}$ to $\frac{\sigma^{2}}{\lambda_{k_{0}}^{1}}=\frac{\sigma^{2}}{\lambda_{k_{0}}}+1$. In the subsequent time slots, the remaining $Q-1$ ice blocks are filled onto the channels locating at the deepest base surfaces or ice levels, indexed by $k_{t}=\arg\min_{k}\{\frac{\sigma^{2}}{\lambda_{k}^{t}}\}$, one by one. The corresponding eigenvectors ${\mathbf{u}}_{k_{t}}$ are used for pilot transmission, and the ice levels increase according to Lemma 2. Consequently, the final pilot reuse frequencies $\{n_{k}\}$ are determined by the number of ice blocks on top of each channel, and the ice levels are given by $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}=\{n_{k}+\frac{\sigma^{2}}{\lambda_{k}}\}$.
The comparison between the ice-filling and water-filling algorithms in Fig. 4 indicates that the continuous powers $\{p_{k}\}$ are quantized into the discrete pilot reuse frequencies $\{n_{k}\}$ by the ice-filling algorithm. This quantization process naturally results in the non-uniform ice levels $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}$, as opposed to the uniform water level $\beta$ used in the water-filling principle. To further emphasize the quantization nature of ice-filling, we prove Theorem 2 to upper-bound the quantization error between $p_{k}$ and $n_{k}$.
@@
$\frac{|n_{k}-p_{k}|}{p_{k}}\leq\frac{1}{p_{k}}\overset{Q\rightarrow+\infty}{=}0.$ (19)
@@ IV-D Majorization Minimization for Phase-Only Controllable Combiner
For dense antenna arrays which can only control their phase shifts, the amplitude of each element of $\bf W$ is fixed as a constant. Considering the normalized power constraint, the observation vector at the $(t+1)$-th timeslot can be obtained by
@@ IV-E Kernel Selection in Non-Ideal Case
In practical communication scenarios, obtaining the perfect prior kernel, ${\bf\Sigma}_{\bf h}$, in the initialization states of Algorithms 2 and 3 may pose a challenge. To deal with this issue, we suggest two methods to obtain imperfect prior kernels.
@@ IV-E2 Artificial kernels
  • Exponential kernel: The exponential kernel ${\bm{\Sigma}}_{\rm exp}$, is the most popular selection in Bayesian estimation. Let ${{\bf{m}}_{1}}=\left[{-\frac{{{M_{1}}-1}}{2},-\frac{{{M_{1}}-3}}{2},\cdots,\frac{{{M_{1}}-1}}{2}}\right]^{T}$ and ${{\bf{m}}_{2}}=\left[{-\frac{{{M_{2}}-1}}{2},-\frac{{{M_{2}}-3}}{2},\cdots,\frac{{{M_{2}}-1}}{2}}\right]^{T}$. Then, the exponential kernels for the two dimensions of UPA, ${{\bf{\Sigma}}_{{\rm exp},1}}\in{\mathbb{C}}^{M_{1}\times M_{1}}$ and ${{\bf{\Sigma}}_{{\rm exp},2}}\in{\mathbb{C}}^{M_{2}\times M_{2}}$, can be respectively expressed as
    ${{\bf{\Sigma}}_{{\rm exp},1}}\!=\!\exp\left({\!-{\eta_{1}^{2}}{\frac{4\pi^{2}d^{2}}{\lambda^{2}}}{{\left|{{{\bf{1}}^{T}_{{M_{1}}}}\otimes{{\bf{m}}_{1}}\!-\!{\bf{m}}_{1}^{T}\otimes{{\bf{1}}_{{M_{1}}}}}\right|}^{\odot 2}}}\right),$ (25)
    ${{\bf{\Sigma}}_{{\rm exp},2}}\!=\!\exp\left({\!-{\eta_{1}^{2}}{\frac{4\pi^{2}d^{2}}{\lambda^{2}}}{{\left|{{{\bf{1}}^{T}_{{M_{2}}}}\otimes{{\bf{m}}_{2}}\!-\!{\bf{m}}_{2}^{T}\otimes{{\bf{1}}_{{M_{2}}}}}\right|}^{\odot 2}}}\right),$ (26)
@@
    ${\bm{\Sigma}}_{\rm exp}={{\bf{\Sigma}}_{{\rm exp},1}}\otimes{{\bf{\Sigma}}_{{\rm exp},2}}.$ (27)
@@
    ${{\bf{\Sigma}}_{{\rm{bes}},1}}$ $={J_{0}}\left({\eta_{2}\frac{2\pi d}{\lambda}\left|{{\bf{1}}_{{M_{1}}}^{T}\otimes{{\bf{m}}_{1}}-{\bf{m}}_{1}^{T}\otimes{{\bf{1}}_{{M_{1}}}}}\right|}\right),$ (28)
    ${{\bf{\Sigma}}_{{\rm{bes}},2}}$ $={J_{0}}\left({\eta_{2}\frac{2\pi d}{\lambda}\left|{{\bf{1}}_{{M_{2}}}^{T}\otimes{{\bf{m}}_{2}}-{\bf{m}}_{2}^{T}\otimes{{\bf{1}}_{{M_{2}}}}}\right|}\right),$ (29)
@@
    ${{\bf{\Sigma}}_{{\rm{bes}}}}={{\bf{\Sigma}}_{{\rm{bes}},1}}\otimes{{\bf{\Sigma}}_{{\rm{bes}},2}}.$ (30)
@@ V Performance Analysis
To evaluate the performance of the proposed MIM-based channel estimators, this section provides comprehensive MSE analyses under different algorithmic conditions.
@@ V-A Estimation Accuracy Under Perfect Kernel
See Appendix E. ∎
We can gain two insights from Lemma 3 and Lemma 4. Firstly, the MSEs achieved by Algorithm 1 and Algorithm 2 have identical asymptotic expressions, both of which decay at the rate of $Q^{-1}$. The only difference is owing to the quantization of the continuous powers $\{p_{k}\}$ in (32) to the pilot reuse frequencies $\{n_{k}\}$ in (33). Next, the MSEs decay quadratically as the rank of the channel kernel, $K$, decreases. This finding demonstrates the superiority of DASs in channel estimation. A denser antenna deployment increases the correlation between inter-antenna channels, decreases the rank of the channel kernel, and thereby improves the channel estimation accuracy.
@@
See Appendix G. ∎
It is evident from Lemmas 3-5 that meticulously designed observation matrices can scale down the asymptotic MSE from $\mathcal{O}(KMQ^{-1})$ to $\mathcal{O}(K^{2}Q^{-1})$. This improvement is attributed to the fact that the observation matrices designed by water-filling and ice-filling approaches are tightly aligned with the kernel’s eigenspace of dimension $K$. In contrast, the random observation matrix wastes a large amount of power in the kernel’s null space of dimension $M-K$, which contains no useful information about the channel. This finding demonstrates the significant role played by the observation matrix to enhance channel estimation accuracy, particularly in DASs where $M\gg K$.
@@ V-B Estimation Accuracy Under Imperfect Kernel
See Appendix H. ∎
Lemma 6 indicates that, the estimation error caused by the non-ideal input cannot be neglected. For example, if the input kernel $\bm{\Sigma}$ behaves far from the real kernel ${\bm{\Sigma}}_{\bf h}$, the value of the MSE may become significantly large due to the existence of $\bm{\Pi}$ in (36). In this case, one may be concerned about the optimality and the accessibility of the proposed estimator when $\bf\Sigma$ can be arbitrarily selected. To answer the question, we derive the following corollary.
Corollary 1
The lower-bound of the achievable MSE ${\hat{\delta}}$ is exactly the trace of posterior covariance ${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}$ under perfect kernel, i.e., $\mathop{\min}\nolimits_{{\bm{\Sigma}}}~{}{\hat{\delta}}=\mathsf{Tr}\left({{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}\right)$, which is achieved if and only if the input kernel satisfies ${\bf{\Sigma}}={{\bf{\Sigma}}_{\bf{h}}}$.
Proof:
See Appendix I. ∎
Coinciding with the well-known linear estimation theory, the above corollary is proved from the other perspective, and suggests that the real kernel is actually the best prior knowledge for Bayesian regression. It also reveals the equivalence between a kernel-based Bayesian reconstructor and a classical MMSE estimator in DAS channel estimation.
To obtain more insightful results, we use the statistical kernel, ${\bf\hat{\Sigma}}_{{\mathbf{h}}}$ in (24), as an example to analyze the achievable MSEs. To ease the derivation, we first prove a useful expression of the MSE in Lemma 7.
@@ V-C Computational Complexity Analysis
For the observation matrix design, the complexity of Algorithm 1 is mainly caused by the eigenvalue decomposition and matrix reconstruction, which is ${\cal O}(M^{3}+MKQ)$. The complexity of Algorithm 2 is mainly caused by the eigenvalue decomposition and the principal eigenvalue selection, which is ${\cal O}(M^{3}+QK)$. The complexity of Algorithm 3 mainly depends on the search for the largest eigenvalue and the iterations required for the convergence of MM, which is ${\cal O}(M^{3}+I_{o}M^{2}Q)$ wherein $I_{o}$ is the number of iterations. and then saved in the memory medium in advance.
@@ VI-A Simulation Setup and Benchmarks
To account for a practical scenario, the standard 3GPP TR 38.901 channel model is used for simulations, whose key parameters are given in Table I. Otherwise specifically specified, we consider a UPA. The number of antennas is set to $M=128$, and the numbers of horizontal antennas and vertical antennas are set to $M_{x}=16$ and $M_{y}=8$, respectively. Assume that the antenna spacing is set to $\frac{\lambda}{8}$. The signal-to-noise ratio (SNR) is defined as ${\rm SNR}=\frac{{\mathsf{E}}\left(\|{\bf h}\|^{2}\right)}{\sigma^{2}}$, of which the default value is set to $10$ dB. Let $\hat{\bf h}$ denote the estimated value of channel $\bf h$. The performance is evaluated by the normalized mean square error (NMSE), which is defined as ${\rm NMSE}={\mathsf{E}}\left(\frac{\|{\bf h}-\hat{\bf h}\|^{2}}{\|{\bf h}\|^{2}}\right)$. Besides, 3000 Monte-Carlo experiments are carried out to plot each figure.
To demonstrate the effectiveness of the proposed ice-filling and MM enabled channel estimators, seven channel estimation schemes are considered for comparison:
  • LS: The classic LS method is applicable when the pilot length is no less than the number of antennas. To implement it, the observation matrix, ${\mathbf{W}}$, is set to a DFT matrix.
  • MMSE: Setting the observation matrix as the DFT matrix, the classic MMSE method is harnessed to recover the channel by using the Bayesian regression in (5).
@@
  • Optimal water-filling: The water-filling scheme in Algorithm 1 is employed to design the ideal observation matrix ${\bf W}$. Based on ${\bf W}$, the Bayesian regression in (5) is performed, which serves as the lower bound of NMSE.
@@
  • Proposed ice-filling: The ice-filling scheme in Algorithm 2 is employed to design the observation matrix ${\bf W}$. Then, ${\bf W}$ is used for Bayesian regression in (5).
  • Proposed MM: The MM-based scheme in Algorithm 3 is employed to design the observation matrix ${\bf W}$. Then, ${\bf W}$ is used for Bayesian regression in (5).

### v2

@@ IV-C Ice-Filling Versus Water-Filling
In this subsection, we put forward a more insightful interpretation to the ice-filling algorithm that establishes its connection with the water-filling principle. Generally speaking, the proposed ice-filling algorithm can be viewed as a quantization of the water-filling principle. Specifically, it is clear that the observation matrices generated by both the water-filling and the ice-filling algorithms fall within the eigenspace of ${\bm{\Sigma}}_{\bf h}$. The major distinction is attributed to their “power allocation strategies”. The water-filling principle allocates $Q$ units of power to the $K$ eigenvectors using (7). On the other hand, the ice-filling algorithm transforms this power-allocation process into an eigenvector-assignment process, where the $K$ eigenvectors are repeatably assigned to the observation vectors in $Q$ timeslots.
To be more precise, recall that $p_{k}$ in (7) denotes the optimal power allocated to the $k$-th eigenvector, satisfying $\sum_{k=1}^{K}p_{k}=Q$. For a clear comparison, we introduce the definition of “pilot reuse frequency” as follows.
@@
Comparing the definitions of $p_{k}$ and $n_{k}$, we can intuitively interpret the $n_{k}$-timeslot reuse of the observation vector ${\mathbf{u}}_{k}$ for pilot transmission as allocating $n_{k}$ units of power to ${\mathbf{u}}_{k}$. More importantly, Definition 1 enables us to derive the analytical expressions of the eigenvalues $\{\lambda_{k}^{t}\}$ in Algorithm 1. Note that the eigenvalue update rule of the ice-filling algorithm can be rewritten as a recursive formula:
@@
where $\lambda_{k_{t}}^{t}$ is the largest eigenvalue from $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$. This recursive formula indicates that whenever the $k$-th eigenvector is selected as an observation vector, the value of $\frac{\sigma^{2}}{\lambda_{k}^{t}}$ increases by 1. By further considering the definition of the pilot reuse frequency $n_{k}^{t}$, that the $k$-th eigenvector is selected by $n_{k}^{t}$ times during timeslots $1\sim t$, we can naturally obtain the relationship between $n_{k}^{t}$ and $\frac{\sigma^{2}}{\lambda_{k}^{t}}$:
@@
The comparison between equations (16) and (7) allows us to interpret the ice-filling algorithm as a quantization of the water-filling algorithm, as elaborated below.
Interpretation of ice-filling: As depicted in Fig. 3, the water-filling principle allocates $Q$ units of water (power) to a vessel with $K$ channels, each having a unique base level $\frac{\sigma^{2}}{\lambda_{k}}$, $k\in\{1,2,\cdots,K\}$. By controlling the uniform water level $\beta$, the optimal power $p_{k}$ is determined by the gap between the base level and the water level. In contrast, the ice-filling algorithm transforms the total $Q$ units of water into $Q$ ice blocks, each containing one unit of power and being used for one pilot transmission. Our algorithm starts from an empty vessel and fills one ice block onto the channel having the deepest base surface $\frac{\sigma^{2}}{\lambda_{k_{0}}}$. This operation is equivalent to finding the largest eigenvalue $\lambda_{k_{0}}$ from $\{\lambda_{1},\cdots,\lambda_{K}\}$. Then, the eigenvector ${\mathbf{u}}_{k_{0}}$ is assigned to the first observation vector ${\mathbf{w}}_{1}$, and ice level of this channel increases from $\frac{\sigma^{2}}{\lambda_{k_{0}}}$ to $\frac{\sigma^{2}}{\lambda_{k_{0}}^{1}}=\frac{\sigma^{2}}{\lambda_{k_{0}}}+1$. In the subsequent time slots, the remaining $Q-1$ ice blocks are filled onto the channels locating at the deepest base surfaces or ice levels, indexed by $k_{t}=\arg\min_{k}\{\frac{\sigma^{2}}{\lambda_{k}^{t}}\}$, one by one. The corresponding eigenvectors ${\mathbf{u}}_{k_{t}}$ are used for pilot transmission, and the ice levels increase according to (16). Consequently, the final pilot reuse frequencies $\{n_{k}\}$ are determined by the number of ice blocks on top of each channel, and the ice levels are given by $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}=\{n_{k}+\frac{\sigma^{2}}{\lambda_{k}}\}$.
The comparison between the ice-filling and water-filling in Fig. 3 indicates that the continuous powers $\{p_{k}\}$ are quantized into the discrete pilot reuse frequencies $\{n_{k}\}$. This quantization results in the non-uniform ice levels $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}$, as opposed to the uniform water level $\beta$. To see the quantization nature of ice-filling more clearly, we prove Theorem 2 to upper bound the quantization error between $p_{k}$ and $n_{k}$.
@@
$\frac{|n_{k}-p_{k}|}{p_{k}}<\frac{1}{p_{k}}\overset{Q\rightarrow+\infty}{=}0.$ (18)
@@ IV-D Majorization Minimization for Phase-Only Controllable Combiner
Turn now to the phase-only controllable analog combiners where the amplitude of each element of $\bf W$ is fixed as a constant. The observation vector at the $(t+1)$-th timeslot can be obtained by
@@ IV-E Kernel Selection in Non-Ideal Case
In practical communication scenarios, obtaining the perfect prior kernel, ${\bf\Sigma}_{\bf h}$, in the initialization states of Algorithms 1 and 2 may pose a challenge. To deal with this issue, we suggest two methods to obtain imperfect prior kernels.
@@ IV-E2 Artificial kernels
  • Exponential kernel: The exponential kernel ${\bm{\Sigma}}_{\rm exp}$, is the most popular selection in Bayesian estimation. Let ${{\bf{m}}_{x}}=\left[{-\frac{{{M_{x}}-1}}{2},-\frac{{{M_{x}}-3}}{2},\cdots,\frac{{{M_{x}}-1}}{2}}\right]^{T}$ and ${{\bf{m}}_{y}}=\left[{-\frac{{{M_{y}}-1}}{2},-\frac{{{M_{y}}-3}}{2},\cdots,\frac{{{M_{y}}-1}}{2}}\right]^{T}$. Then, the exponential kernels for the two dimensions of UPA, ${{\bf{\Sigma}}_{{\rm exp},x}}\in{\mathbb{C}}^{M_{x}\times M_{x}}$ and ${{\bf{\Sigma}}_{{\rm exp},y}}\in{\mathbb{C}}^{M_{y}\times M_{y}}$, can be respectively expressed as
    ${{\bf{\Sigma}}_{{\rm exp},x}}\!=\!\exp\left({\!-{\eta_{1}^{2}}{\frac{4\pi^{2}d^{2}}{\lambda^{2}}}{{\left|{{{\bf{1}}^{T}_{{M_{x}}}}\otimes{{\bf{m}}_{x}}\!-\!{\bf{m}}_{x}^{T}\otimes{{\bf{1}}_{{M_{x}}}}}\right|}^{\odot 2}}}\right),$ (24)
    ${{\bf{\Sigma}}_{{\rm exp},y}}\!=\!\exp\left({\!-{\eta_{1}^{2}}{\frac{4\pi^{2}d^{2}}{\lambda^{2}}}{{\left|{{{\bf{1}}^{T}_{{M_{y}}}}\otimes{{\bf{m}}_{y}}\!-\!{\bf{m}}_{y}^{T}\otimes{{\bf{1}}_{{M_{y}}}}}\right|}^{\odot 2}}}\right),$ (25)
@@
    ${\bm{\Sigma}}_{\rm exp}={{\bf{\Sigma}}_{{\rm exp},x}}\otimes{{\bf{\Sigma}}_{{\rm exp},y}}.$ (26)
@@
    ${{\bf{\Sigma}}_{{\rm{bes}},x}}$ $={J_{0}}\left({\eta_{2}\frac{2\pi d}{\lambda}\left|{{\bf{1}}_{{M_{x}}}^{T}\otimes{{\bf{m}}_{x}}-{\bf{m}}_{x}^{T}\otimes{{\bf{1}}_{{M_{x}}}}}\right|}\right),$ (27)
    ${{\bf{\Sigma}}_{{\rm{bes}},y}}$ $={J_{0}}\left({\eta_{2}\frac{2\pi d}{\lambda}\left|{{\bf{1}}_{{M_{y}}}^{T}\otimes{{\bf{m}}_{y}}-{\bf{m}}_{y}^{T}\otimes{{\bf{1}}_{{M_{y}}}}}\right|}\right),$ (28)
@@
    ${{\bf{\Sigma}}_{{\rm{bes}}}}={{\bf{\Sigma}}_{{\rm{bes}},x}}\otimes{{\bf{\Sigma}}_{{\rm{bes}},y}}.$ (29)
@@ V Performance Analysis
To evaluate the performance of the proposed MIM-based observation matrix design, this section provides comprehensive MSE analyses under different algorithmic conditions.
@@ V-A Estimation Accuracy Under Perfect Kernel
See Appendix D. ∎
We can gain two insights from Lemma 2 and Lemma 3. Firstly, the MSEs achieved by the water-filling and ice-filling methods have identical asymptotic expressions, both of which decay at the rate of $Q^{-1}$. The only difference is owing to the quantization of the continuous powers $\{p_{k}\}$ in (31) to the pilot reuse frequencies $\{n_{k}\}$ in (32). Next, the MSEs decay quadratically as the rank of the channel kernel, $K$, decreases. This finding demonstrates the superiority of DASs in channel estimation. A denser antenna deployment increases the correlation between inter-antenna channels, decreases the rank of the channel kernel, and thereby improves the channel estimation accuracy.
@@
See Appendix F. ∎
It is evident from Lemmas 2-4 that meticulously designed observation matrices can scale down the asymptotic MSE from $\mathcal{O}(KMQ^{-1})$ to $\mathcal{O}(K^{2}Q^{-1})$. This improvement is attributed to the fact that the observation matrices designed by water-filling and ice-filling approaches are tightly aligned with the kernel’s eigenspace of dimension $K$. In contrast, the random observation matrix wastes a large amount of power in the kernel’s null space of dimension $M-K$, which contains no useful information about the channel. This finding demonstrates the significant role played by the observation matrix to enhance channel estimation accuracy, particularly in DASs where $M\gg K$.
@@ V-B Estimation Accuracy Under Imperfect Kernel
See Appendix G. ∎
Lemma 5 indicates that, the estimation error caused by the non-ideal input cannot be neglected. For example, if the input kernel $\bm{\Sigma}$ behaves far from the real kernel ${\bm{\Sigma}}_{\bf h}$, the value of the MSE may become significantly large due to the existence of $\bm{\Pi}$ in (35). To obtain more insightful results, we use the statistical kernel, ${\bf\hat{\Sigma}}_{{\mathbf{h}}}$ in (23), as an example to analyze the achievable MSEs. To ease the derivation, we first prove a useful expression of the MSE in Lemma 6.
@@ V-C Computational Complexity Analysis
For the observation matrix design, the complexity of Algorithm 1 is mainly caused by the eigenvalue decomposition and the principal eigenvalue selection, which is ${\cal O}(M^{3}+QK)$. The complexity of Algorithm 2 mainly depends on the search for the largest eigenvalue and the iterations required for the convergence of MM, which is ${\cal O}(M^{3}+I_{o}M^{2}Q)$ wherein $I_{o}$ is the number of iterations. and then saved in the memory medium in advance.
@@ VI-A Simulation Setup and Benchmarks
To account for a practical scenario, the standard 3GPP TR 38.901 channel model is used for simulations, whose key parameters are given in Table I. Otherwise specifically specified, we consider a UPA. The number of antennas is set to $M=128$, and the numbers of horizontal antennas and vertical antennas are set to $M_{x}=16$ and $M_{y}=8$, respectively. Assume that the antenna spacing is set to $\frac{\lambda}{8}$. The signal-to-noise ratio (SNR) is defined as ${\rm SNR}=\frac{{\mathsf{E}}\left(\|{\bf h}\|^{2}\right)}{\sigma^{2}}$, of which the default value is set to $10$ dB. The performance is evaluated by the normalized mean square error (NMSE), which is defined as ${\rm NMSE}={\mathsf{E}}\left(\frac{\|{\bf h}-\hat{\bf h}\|^{2}}{\|{\bf h}\|^{2}}\right)$. Besides, 3000 Monte-Carlo experiments are carried out to plot each figure.
To demonstrate the effectiveness of the proposed ice-filling and MM enabled channel estimators, five benchmark schemes are considered for comparison:
  • LS: The classic LS method is applicable when the pilot length is no less than the number of antennas. To implement it, the pilot length is fixed as $Q=M$ and the observation matrix, ${\mathbf{W}}$, is set to a DFT matrix.
  • MMSE: The MMSE benchmark utilizes the same DFT observation matrix as the LS scheme. It then leverages the Bayesian regression in (4) to recover the channel.
@@
  • Ideal water-filling: The water-filling scheme is employed to design the ideal observation matrix ${\bf W}$. The noise vector, ${\mathbf{z}}$, in (2) is ideally assumed to be independent of the observation matrix ${\mathbf{W}}$, and its distribution is fixed as $\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$.